1.A 25 kg weather balloon experiences a buoyant force of 710 N of upwards. Determine how long it takes for the balloon to reach a height of 1.0 km if it is released from the ground. (Hint: draw a FBD for the weather balloon on a separate piece of paper as your rough work)

2.Former Toronto Blue Jays center fielder Kevin Pillar is stealing 2nd base. He begins his slide 5.2 m from the base while running at his top speed of 4.5 m/s. If Kevin comes to a complete stop exactly on 2nd base, determine the force of friction acting on Kevin if his mass is 93 kg.

3.


1
Expert's answer
2021-07-16T08:51:51-0400

1)

"ma=F_b-mg\\25a=710-25(9.8)\\a=18.6\frac{m}{s^2}\\0.5(18.6)t^2=1000\\t=10.4\ s"

2)


"mv^2=2F_fd\\(93)(4.5)^2=2(5.2)F_f\\F_f=180\ N"